# ST-FlowNet: An Efficient Spiking Neural Network for Event-Based Optical Flow Estimation

### v1

@@ 3.3.2 Bio-information-fused training
To address these challenges, we propose a bio-information-fused training strategy (BISNN). This approach incorporates two key operations: (1) Cross-model initialization: The SNN models, which share an identical architecture, are initialized with the pre-trained weights $\left\{w^{l}\right\}_{l=0}^{L}$ of the ANN models. This facilitates efficient knowledge transfer between models; and (2) Parameter-free optimization: A supervised retraining procedure utilizing the STBP (Wu et al., 2018) method is employed to optimize the SNN models, thereby circumventing the need for complex biological parameter filtering processes. According to the chain rule, the mathematical formulation of the loss function’s derivatives with respect to the learnable parameters $\left\{w^{l}\right\}_{l=0}^{L}$ can be expressed as follows:
@@
Table 2: Comparison of state-of-the-art models across different datasets. The ST-FlowNet model trained using the ANN, A2S, and BISNN methods are denoted as ST-FlowNet1, ST-FlowNet2, and ST-FlowNet3, respectively. For the MVSEC dataset, the AEE results for each scenario are presented. For the ECD and HQF datasets, the average FWL and RSAT results across all scenarios are reported. The best results are highlighted in bold. The symbol $\downarrow$ ($\uparrow$) indicates that a smaller (larger) value is preferred.
@@ 4.1.1 Optical Flow Datasets
4.1.2 Evaluation Metrics
The ECD and HQF datasets do not include ground truth, and thus we employ two image compensation quality metrics to estimate predicted flows. Specifically, the flow warp loss (FWL) assesses the sharpness of the image of warped events compared with the original event partition, and the variance of the contrast of the event images is reported (T. Stoffregen, 2020; Gallego et al., 2019, pp. 12280–12289). Additionally, we also report the ratio of the squared average timestamps (RSAT), indicating the contrast of $\mathcal{L}_{\mathrm{contrast}}$ between predicted optical flow and baseline null vectors (Hagenaars et al., 2021).
4.1.3 Implementation Details
4.2 Comparison with State-of-the-art Methods
To validate the effectiveness of our proposed ST-FlowNet, we conduct a comprehensive comparative analysis, assessing its performance compared with other state-of-the-art models from both quantitative and qualitative perspectives (Hagenaars et al., 2021; Scheerlinck et al., 2020; Tian and Andrade-Cetto, 2022). The ST-FlowNet model trained using the ANN, A2S, and BISNN methods are denoted as ST-FlowNet1, ST-FlowNet2, and ST-FlowNet3, respectively.
4.2.1 Quantitative Evaluation
We conduct a comparative analysis between ST-FlowNet and other state-of-the-art models on the MVSEC dataset across four representative scenarios: outdoor_day1 (denoted as OD1), indoor_flying1 (IF1), indoor_flying2 (IF2) and indoor_flying3 (IF3), considering both the dt=1 and dt=4 conditions. As shown in Tab. 2, ST-FlowNet demonstrates superior AEE performance compared with other ANN and SNN models across most of scenarios. Notably, the optical flow estimation performance can be improved slightly by leveraging the converted SNN model, as particularly evident in the OD1 and IF2 scenarios. For the sake of comparison, on the ECD and HQF dataset, we present a comparative analysis of state-of-the-art models without distinguishing scenarios. As shown in Tab. 2, ST-FlowNet attains the optimal performance, excelling with respect to both the FWL and RSAT metrics.
@@ 4.3.1 Analysis of the Number of Decoders
We compare the baseline models with varying numbers of decoders to examine the effectiveness of the decoders. Specifically, we increase the number of decoders from one in the baseline model to four in the comparison models. To conduct a thorough comparison, we select four representative scenarios from the ECD and HQF datasets, each denoted as follows: dynamic_6dof (D6), boxes_6dof (B6), poster_6dof (P6), slider_depth (SD), along with boxes (BO), engineering_posters (EP), high_texture_plants (HTP), and reflective_materials (RM). As shown in Tabs. 3 and 4, the performance decreases across all scenarios when more decoders are employed. This suggests that the use of more decoders is not necessarily associated with improved performance. This observation not only alleviates the pressure to increase the model size but also reinforces the advantage of the SNN model in terms of energy consumption.
@@ 4.3.2 Analysis of the Spatio-temporal Augmentation
Building on the baseline model with a single decoder, we integrate ConvGRU1 as the spatio-temporal optical flow augmentation component. Unlike the complete ST-FlowNet, the ConvGRU1 layer receives the upsampled basic predicted optical flow as state information. The detailed results are shown in Tabs. 3 and 4. The AEE results for the MVSEC dataset demonstrate that the ConvGRU augmentation layer enhances optical flow estimation performance in most scenarios, except for OD1. The FWL and RSAT results computed for the ECD and HQF datasets also generally corroborate the improvement, which is more pronounced for the HQF dataset. Furthermore, the ConvGRU layers assist in mitigating the performance loss of converted SNN models. In certain scenarios, SNN models exhibit superior performance compared with ANN models, as observed for the IF2 and IF3 scenarios.
@@ 4.3.3 Analysis of the Spatio-temporal Alignment
Expanding on the previous model, we continue to introduce ConvGRU2 as a spatio-temporal alignment module. This module is designed to project the basic and historical predicted optical flow into a standardized state space. By doing so, it not only enhances the precision of optical flow prediction but also establishes an effective reference state for input data. As illustrated in Tabs. 3 and 4, the optical flow estimation performance further improves in most scenarios. Moreover, in comparison to the spatio-temporal augmentation module, the alignment module demonstrates superior effectiveness.
@@ 5 Discussion and Conclusion
In this study, we present a novel neural network architecture, termed ST-FlowNet, which incorporates enhanced ConvGRU layers, designed to align and augment spatio-temporal information, thereby improving optical flow estimation based on event-driven inputs. Our experimental results validate the effectiveness of the proposed ST-FlowNet, demonstrating its superior performance compared to other state-of-the-art models. Ablation studies further highlight the critical role of the integrated components in our architecture for achieving performance improvements.
To address the challenges in training SNN models for optical flow estimation, we propose two methods: (1) the A2S method, which generates an SNN model from a pre-trained ANN model, and (2) a novel BISNN strategy, aimed at mitigating the complexities associated with the selection of biological parameters. Both methods significantly enhance model performance when compared to the direct training method using STBP. Notably, the BISNN method demonstrates greater robustness across various initialized biological parameters.
The limitations of the ST-FlowNet model are discussed below. First, we observe performance degradation in the converted SNN models under specific conditions (e.g., IF1 in Tab. 2 and HTP in Tab. 4), which may be attributed to theoretical conversion errors inherent in the A2S process (Deng and Gu, 2021). This issue could be addressed through the development of more efficient conversion methods. Second, the spatio-temporal augmentation module exhibits a less robust ability to improve performance, potentially due to modality bias between the input event data and the optical flow state. A well-designed ConvGRU1 module may further improve the performance of the ST-FlowNet model by reducing modality bias, deserving to be examined in our further studies.

### v2

@@ 4.3.2 Bio-information-fused training
To address these challenges, we propose a hybrid bio-information-fused training strategy (BISNN). This approach incorporates two key operations: (1) Cross-model initialization: The SNN models, which share an identical architecture, are initialized with the pre-trained weights $\left\{w^{l}\right\}_{l=0}^{L}$ of the ANN models. This facilitates efficient knowledge transfer between models; and (2) Parameter-free optimization: A supervised retraining procedure utilizing the STBP (Wu et al., 2018) method is employed to optimize the SNN models, thereby circumventing the need for complex biological parameter filtering processes. According to the chain rule, the mathematical formulation of the loss function’s derivatives with respect to the learnable parameters $\left\{w^{l}\right\}_{l=0}^{L}$ can be expressed as follows:
@@
Table 2: Comparison of state-of-the-art models across different datasets. The ST-FlowNet model trained using the ANN, A2S, and BISNN methods are denoted as ST-FlowNet1, ST-FlowNet2, and ST-FlowNet3, respectively. For the MVSEC dataset, the AEE results for each scenario are presented. For the ECD and HQF datasets, the average FWL and RSAT results across all scenarios are reported. The optimal and suboptimal results are highlighted in bold. The symbol $\downarrow$ ($\uparrow$) indicates that a smaller (larger) value is preferred.
@@
Adaptive-SpikeNet (Kosta and Roy, 2023) 0.44 0.79 1.37 1.11
SDformerFlow (Tian and Andrade-Cetto, 2025) 0.69 0.61 0.83 0.76
FSFN (Apolinario and Roy, 2024) 0.50 0.76 1.19 1.00
@@
Table 3: Comparison of models trained or fine-tuned on the MVSEC dataset.
        Model         OD1         IF1         IF2         IF3
        RNN-FireNet-S-FT (Xu et al., 2025b)         1.97 (AEE4)         3.24 (AEE4)         5.48 (AEE4)         4.45 (AEE4)
        STE-FlowNet (Ding et al., 2022)         0.42 (AEE1)         0.57 (AEE1)         0.79 (AEE1)         0.72 (AEE1)
        U-Net-like SNN (Cuadrado et al., 2023)         –         0.58 (AEE1)         0.72 (AEE1)         0.67 (AEE1)
        ST-FlowNet2 (ours)         0.37 (AEE1)         0.50 (AEE1)         0.84 (AEE1)         0.70 (AEE1)
        1.24 (AEE4)         1.86 (AEE4)         2.78 (AEE4)         2.34 (AEE4)
@@ 5.1.1 Optical Flow Datasets
5.1.2 Evaluation Metrics
The ECD and HQF datasets do not include ground truth, and thus we employ two image compensation quality metrics to estimate predicted flows. Specifically, the flow warp loss (FWL) assesses the sharpness of the image of warped events compared with the original event partition, and the variance of the contrast of the event images is reported (T. Stoffregen, 2020). Additionally, we also report the ratio of the squared average timestamps (RSAT), indicating the contrast of $\mathcal{L}_{\mathrm{contrast}}$ between predicted optical flow and baseline null vectors (Hagenaars et al., 2021).
5.1.3 Implementation Details
5.2 Comparison with State-of-the-art Methods
To validate the effectiveness of our proposed ST-FlowNet, we conduct a comprehensive comparative analysis, assessing its performance compared with other state-of-the-art models from both quantitative and qualitative perspectives (Scheerlinck et al., 2020; Tian and Andrade-Cetto, 2022). The ST-FlowNet model trained using the ANN, A2S, and BISNN methods are denoted as ST-FlowNet1, ST-FlowNet2, and ST-FlowNet3, respectively.
5.2.1 Quantitative Evaluation
We conduct a comparative analysis between ST-FlowNet and other state-of-the-art models on the MVSEC dataset across four representative scenarios: outdoor_day1 (denoted as OD1), indoor_flying1 (IF1), indoor_flying2 (IF2) and indoor_flying3 (IF3), considering both the dt=1 and dt=4 conditions. As shown in Tab. 2, ST-FlowNet demonstrates superior AEE performance compared with other ANN and SNN models across most of scenarios. Notably, the optical flow estimation performance can be improved slightly by leveraging the converted SNN model, as particularly evident in the OD1 and IF2 scenarios. We also provide a comparison between our ST-FlowNet2 SNN model and other models that are directly trained or fine-tuned on the MVSEC dataset. As shown in Tab. 3, ST-FlowNet2 still demonstrates competitive performance compared to other models. We believe that training on a larger dataset may further enhance the optical flow estimation performance of our model. For the sake of comparison, on the ECD and HQF dataset, we present a comparative analysis of state-of-the-art models without distinguishing scenarios. As shown in Tab. 2, ST-FlowNet attains the optimal performance, excelling with respect to both the FWL and RSAT metrics.
@@ 5.3.1 Analysis of the Number of Decoders
We compare the baseline models with varying numbers of decoders to examine the effectiveness of the decoders. Specifically, we increase the number of decoders from one in the baseline model to four in the comparison models. To conduct a thorough comparison, we select four representative scenarios from the ECD and HQF datasets, each denoted as follows: dynamic_6dof (D6), boxes_6dof (B6), poster_6dof (P6), slider_depth (SD), along with boxes (BO), engineering_posters (EP), high_texture_plants (HTP), and reflective_materials (RM). As shown in Tabs. 4 and 5, the performance decreases across all scenarios when more decoders are employed. This suggests that the use of more decoders is not necessarily associated with improved performance. This observation not only alleviates the pressure to increase the model size but also reinforces the advantage of the SNN model in terms of energy consumption.
@@ 5.3.2 Analysis of the Spatio-temporal Augmentation
Building on the baseline model with a single decoder, we integrate ConvGRU1 as the spatio-temporal optical flow augmentation component. Unlike the complete ST-FlowNet, the ConvGRU1 layer receives the upsampled basic predicted optical flow as state information. The detailed results are shown in Tabs. 4 and 5. The AEE results for the MVSEC dataset demonstrate that the ConvGRU augmentation layer enhances optical flow estimation performance in most scenarios, except for OD1. The FWL and RSAT results computed for the ECD and HQF datasets also generally corroborate the improvement, which is more pronounced for the HQF dataset. Furthermore, the ConvGRU layers assist in mitigating the performance loss of converted SNN models. In certain scenarios, SNN models exhibit superior performance compared with ANN models, as observed for the IF2 and IF3 scenarios.
@@ 5.3.3 Analysis of the Spatio-temporal Alignment
Expanding on the previous model, we continue to introduce ConvGRU2 as a spatio-temporal alignment module. This module is designed to project the basic and historical predicted optical flow into a standardized state space. By doing so, it not only enhances the precision of optical flow prediction but also establishes an effective reference state for input data. As illustrated in Tabs. 4 and 5, the optical flow estimation performance further improves in most scenarios. Moreover, in comparison to the spatio-temporal augmentation module, the alignment module demonstrates superior effectiveness.
@@ 5.5 Analysis of Energy Consumption
Figure 6: Representative examples of failure cases from our ST-FlowNet models. The regions with failed optical flow estimations are highlighted with red boxes.
5.6 Analysis of Failure Cases
Although our proposed method demonstrates promising performance in optical flow estimation, several challenges remain unresolved. Two representative failure cases are illustrated in Fig. 6. In the first row, the ST-FlowNet models fail to accurately capture a small object, as highlighted by the red box. This observation is consistent with the visualization results in Fig. 3, which indicate a relatively degraded performance in scenarios containing numerous small-scale objects (i.e. D6 and RM). In addition, distinguishing between background noise and high-density event streams presents another significant challenge, as shown in the second row of Fig. 6. This issue, common in event-based vision tasks, may potentially be addressed through the development of more suitable event representation strategies.
@@ 6 Discussion and Conclusion
In this study, we present a novel neural network architecture, termed ST-FlowNet, which incorporates enhanced ConvGRU layers, designed to align and augment spatio-temporal information, thereby improving optical flow estimation based on event-driven inputs. To address the challenges in training SNN models for optical flow estimation, we propose two methods: (1) the A2S method, which generates an SNN model from a pre-trained ANN model, and (2) a novel BISNN strategy, aimed at mitigating the complexities associated with the selection of biological parameters.
Overall, our work demonstrates a notable level of superiority. (1) Our experimental results across a variety of representative scenarios validate the effectiveness of the proposed ST-FlowNet, which outperforms current state-of-the-art optical flow estimation models. In particular, for the challenging boundary regions characterized by motion blur or sparse features (commonly introduced during event representation), ST-FlowNet exhibits a robust capability for capturing scene information. (2) Ablation studies further highlight the critical roles of the integrated ConvGRU modules in spatio-temporal augmentation and alignment, establishing a promising tool for future model design. Meanwhile, increasing the number of decoder layers is shown to be non-essential in semi-pyramidal architectural models, thereby contributing to a lightweight and energy-efficient design. (3) A comparative analysis of three training paradigms demonstrates that indirect A2S conversion and hybrid BISNN methods can generate superior SNN models compared to direct STBP training. Notably, the BISNN method exhibits greater robustness to variations in initialized biological parameters, thereby alleviating the complexity of biological parameter selection.
The limitations and shortcomings of the ST-FlowNet model are discussed below. First, we observe performance degradation in the converted SNN models under specific conditions (e.g., IF1 in Tab. 2 and HTP in Tab. 5), which may be attributed to theoretical conversion errors inherent in the A2S process (Deng and Gu, 2021). This issue could be addressed through the development of more efficient conversion methods. Second, the spatio-temporal augmentation module exhibits a less robust ability to improve performance, potentially due to modality bias between the input event data and the optical flow state. A well-designed ConvGRU1 module may further improve the performance of the ST-FlowNet model by reducing modality bias, deserving to be examined in our further studies. Third, analysis of the failure cases reveals that the ST-FlowNet models exhibit limited sensitivity to small objects, background noise, and high-intensity event streams. Addressing these challenges represents a potential avenue for further enhancing model performance.